If barium is produced by electrolyzing molten barium oxide, how long will it take for 4.8 g of Ba to plate out if 7.5 amps run through the cell? Report your answer in minutes and leave a space between the number and the unit, for example 70 min.
Answer #1

Answer:

Step 1: Explanation

The mass of the the substance deposited during electrolysis can be calculated by using faraday's law of electrolysis

m/E = It / F

Where,  m is the mass of the substance deposited.  

E is equivalent mass of the substance deposited

Equivalent mass(E)= Molar mass / n

Where, n is number of electrons involved

I is current passed in amperes = 7.5 Amperes (given)

t is time for which current is passed

F=Faraday constant , which is equal to 96500 C/mol

Molar mass of Ba = 137.327 g/mol

Step 2: Calculation of mass

The half cell reaction

Ba+2(aq) + 2e- -----> Ba(s)

since Ba+2 is converted into Ba which means n value is 2

So, equivalent mass(E) = 137.327 g/mol divide by 2 = 68.6635 g/mol

substituting the values in above equation

m/E = It / F

time taken(t) = (4.8 g × 96500 C/mol ) / (7.5 amp × 68.6635 g/mol ) = 899.46 s

So, time taken in minute = 899.46 s × (1 min / 60 s ) = 15 min